How do I solve this equation and check for extraneous solutions?
3 sqrt w + 1 = 6

Found 2 solutions by unlockmath, stanbon:
Answer by unlockmath(1688)

Hello,
Let's first divide by 3 to get:
sqrt w + 1 = 2 square this equation to get:
w + 1 = 4 Subtract 1 from both sides to get:
w=3
Make sense?
RJ
www.math-unlock.om

Answer by stanbon(75887)

How do I solve this equation and check for extraneous solutions?
3 sqrt(w + 1) = 6
----
sqrt(w+1) = 2
w+1 = 4
w = 3
-----
Checking:
3*sqrt(4) = 6
3*2 = 6
-----
Solution: w = 3
==================
Cheers,
Stan H.